OpenStudy (anonymous):

solve for X 1/4x+1/2=-4(5/6X+2)

OpenStudy (anonymous):

1/4 x+1/2= -20/6 x-2 get the x's on one side and the whole #'s on the other and solve for x

OpenStudy (anonymous):

but how did you got that

OpenStudy (anonymous):

i got -30/43

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }x +\frac{ 1 }{ 2}=\frac{ -20 }{ 6 } x -2\] \[\frac{ 1 }{ 4 }x+\frac{ 20 }{ 6 }x=-2-\frac{ 1 }{ 2 }\] \[\frac{ (1*3)+(20*2) }{ 12 }=-\frac{ (2*2)-1 }{ 2 }\] \[\frac{ 43 }{ 12 }x=-\frac{ 5 }{ 2 }\] \[x=\frac{ -5 }{ 2 }\times \frac{ 12 }{ 43 }\] \[x=\frac{ -30 }{ 43 }\]

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

thanks

OpenStudy (adunb8):

your right!